y of Physics 203-Fall 2016 Physics2039-Chapter9 Problem 9.48 Problem 9.48 At depths of 2000 m in the sea, the pressure is about 200 times | atmospheric pressure (1 atm = 1.0 × 105 N/m2). The bulk atmospheric pressure(lam,the pressure is about 200times modulus for cast iron is 9.0 × 1010 N/m2 me.
Answer #1

B = -P / (\Delta V/V)

\Delta V/ V = P / B

=200 x 1 x 105 / 9 x 1010

   = 0.2222 x 10-3

% \Delta V/ V  = 0.2222 x 10-1

   = 0.022